OpenStudy (anonymous):

Does the integral converge or diverge? If it converges, find its value. The integral from -3 to 3 1/√ (9-x^2) dx

OpenStudy (anonymous):

\[\int\limits_{-3}^{3?} 1/ \sqrt{9-x^2} dx\]

hartnn (hartnn):

16th in your list

OpenStudy (anonymous):

this is a new section, should of said that. lol were not using this list on this one.

hartnn (hartnn):

lol, but 16th is a standard integral! anyways you can then plug in x= 3 sin y

OpenStudy (anonymous):

i know we should start out by saying lim t -> some number then solve normaly

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

we couldn't use 3 in the denominator

OpenStudy (anonymous):

1/3cosy *3cosy ?

hartnn (hartnn):

oh, for the convergent part, yes you need to show \( \large \lim \limits_{t \rightarrow -3} \int \limits_t^3 ...= finite \quad and \\ \large \lim \limits_{t \rightarrow 3} \int \limits_{-3}^{t} ...= finite\)

hartnn (hartnn):

yup, so, that =1 and its integral = y and y= sin^{-1} x ^_^

hartnn (hartnn):

sorry, here , y= sin^{-1} (x/3) because x= 3 sin y

hartnn (hartnn):

doubts ?

OpenStudy (anonymous):

wouldnt the top and bottom cancel each other out?

hartnn (hartnn):

yeah, the top and bottom cancels out and gives you 1 !!

hartnn (hartnn):

integral of 1 dy is just y+c

OpenStudy (anonymous):

they are trying to be tricky aren't they! lol

OpenStudy (anonymous):

so it would converge since its a number?

hartnn (hartnn):

absolutely, finite number means convergent :)

hartnn (hartnn):

you solved convergent problems before, right ? and know the steps.... ? could find those 2 limits ?

OpenStudy (anonymous):

so how would you do the final subtraction if the answer is just 1.. or y

OpenStudy (anonymous):

cant plug in 3 there?

hartnn (hartnn):

ok, we have \(\int 1dy =y+c \\ y= \sin^{-1}(x/3)\) why not ? sin^{-1} 1 and sin^{-1} -1 are finite numbers (infact angles)!

OpenStudy (anonymous):

ohhh i see what your saying

hartnn (hartnn):

let me give you steps : \(\large \lim \limits_{t \rightarrow -3} \int \limits_t^3 \dfrac{1}{\sqrt{9-x^2}}dx = \lim \limits_{t \rightarrow -3} \int \limits_{\sin^{-1}{(t/3)}}^{\sin^{-1}1}1dy \\ \large =\lim \limits_{t \rightarrow -3} y|_{{\sin^{-1}{(t/3)}}}^{\sin^{-1}1}=\lim \limits_{t \rightarrow -3}{\sin^{-1}{(t/3)}}-{\sin^{-1}1} \\ \large ={\sin^{-1}{(-3/3)}}-{\sin^{-1}1}=....\)

OpenStudy (anonymous):

thanks a lot, that really helped out! its funny how someone on the internet can do a better job explaining then my teacher. lol

hartnn (hartnn):

lol, i get that alot! welcome ^_^